OpenStudy (anonymous):

Which of the following has the greatest mass, in grams? 1 atom of gold 1.0 mol of mercury atoms 2.0 mole of chlorine atoms 25 atoms of mercury

OpenStudy (anonymous):

@Abhisar

OpenStudy (micahm):

first thing first what do you think it is

OpenStudy (anonymous):

(oh thank god finally some help!!) and honestly i have no idea, but if I where to pick one id have to say 25 atoms of mercury(?) idk it sounds like it has the most....... @micahm

OpenStudy (abb0t):

Are you just saying that because it has the largest number, or did you just take a guess? Remember, you can use avogadro's number to convert from atoms to moles, and then the molar mass to get grams of whatever atom you have.

OpenStudy (abb0t):

\(\sf 25~atoms \times \frac{1~mol}{6.022 \times 10^{23}~atoms}(\frac{grams}{mol})_{of~ mercury}\)

OpenStudy (micahm):

@abb0t so then it not the highest but _____________ or the smallest ?

OpenStudy (micahm):

@abb0t

OpenStudy (micahm):

@Secret-Ninja @KamiBug @abb0t

OpenStudy (secret-ninja):

I have no clue >.< Sorry.

OpenStudy (micahm):

some people would chose 25 but accordingly to @abb0t it wrong

OpenStudy (micahm):

@pooja195

pooja195 (pooja195):

http://openstudy.com/study#/updates/53bc4a4ae4b0f3a650f86e0d try that :)

OpenStudy (micahm):

if you can't find the answer go with you first guess

OpenStudy (anonymous):

Don't guess. Rule out the answers with atoms, because they will have incredibly small masses. It is one of the two choices that are given in moles. For each one, convert from moles to mass to be able to determine which mass is higher.

OpenStudy (anonymous):

@pooja195 thanks !!

OpenStudy (abb0t):

Why the heck did you tag me more than once? Clearly, I was not here.